OpenStudy (anonymous):

16^-3/4 evaluate using radical form

hartnn (hartnn):

is it \(16^{-3/4}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you write down the stepss please, then I will understand better, thank you

hartnn (hartnn):

ok,\(16^{(-3/4)}=(16^{1/4})^{-3}\) \[(\sqrt[4]{16})^{-3}=2^{(-3)}=1/2^{3}=1/8\] got it?

hartnn (hartnn):

did u get that?

OpenStudy (anonymous):

where did you get the two from ?

hartnn (hartnn):

ok, \((\sqrt[4]{16})=\sqrt[4]{2*2*2*2}=\sqrt[4]{2^4}=2\) ok?

OpenStudy (anonymous):

okay

hartnn (hartnn):

still doubts with anything?or clear?

OpenStudy (anonymous):

um I think I understand a bit, thank you :)

hartnn (hartnn):

welcome :)